5.3*10^8 g to mama grams . Express your answer to two significant figures and include the appropriate units.


1
Expert's answer
2021-02-04T06:55:26-0500

5.3 × 108 g

1 g = 1000mg

5.3 × 108 g = 5.3 × 108 × 1000 mg

= 5.3 × 1011 mg